# AI-driven Wireless Positioning: Fundamentals, Standards, State-of-the-art, and Challenges

### v1

@@ IX Conclusion
In this work, we primarily explores the potential of AI-driven wireless positioning technologies from the perspective of integrating AI with wireless positioning. While the focus is on cellular positioning scenarios, the study also incorporates insights from WiFi, Bluetooth, and UWB positioning to enhance the comprehensiveness of the algorithmic understanding. Specifically, we introduces the foundational knowledge of AI technologies and wireless positioning techniques, followed by a summary of 3GPP standards related to positioning and AI advancements. Subsequently, we also reviews the SOTA research in both AI/ML-assisted positioning and direct AI/ML positioning, as well as datasets commonly used for wireless positioning. Finally, we summarize the challenges and opportunities in AI-driven wireless positioning. With the continued advancement of AI technologies, the integration of AI and positioning is expected to deepen. This review aims to inspire researchers in both industry and academia, contributing to the advancement of this promising field.

### v2

@@ IX Conclusion
In this work, we primarily explore the potential of AI-driven wireless positioning technologies from the perspective of integrating AI with wireless positioning. While the focus is on cellular positioning scenarios, the study also incorporates insights from WiFi, Bluetooth, and UWB positioning to enhance the comprehensiveness of the algorithmic understanding. Specifically, we introduce the foundational knowledge of AI technologies and wireless positioning techniques, followed by a summary of 3GPP standards related to positioning and AI advancements. Subsequently, we also reviews the SOTA research in both AI/ML-assisted positioning and direct AI/ML positioning, as well as datasets commonly used for wireless positioning. Finally, we summarize the challenges and opportunities in AI-driven wireless positioning. With the continued advancement of AI technologies, the integration of AI and positioning is expected to deepen. This review aims to inspire researchers in both industry and academia, contributing to the advancement of this promising field.